# Revisiting institutional punishment in the $N$-person prisoner’s dilemma

### v2

@@ III The replicator equation formulation
$\frac{dx}{dt}$ $=$ $\frac{x}{\Delta f_{\max}}\left[\pi^{C}(x,z)-\bar{\pi}(x,y,z)\right]$ (13)
$\frac{dy}{dt}$ $=$ $\frac{y}{\Delta f_{\max}}\left[\pi^{D}(x,z)-\bar{\pi}(x,y,z)\right]$ (14)
$\frac{dz}{dt}$ $=$ $\frac{z}{\Delta f_{\max}}\left[\pi^{P}(x,z)-\bar{\pi}(x,y,z)\right],$ (15)
where $\bar{\pi}=x\pi^{C}+y\pi^{D}+z\pi^{P}$ is the average payoff of the population and $\Delta f_{\max}$ is given by Eq. (5). Although we could eliminate the factor $1/\Delta f_{\max}$ by a proper rescaling of $t$ and thus obtain the standard form of the replicator equation, keeping it facilitates comparison with the finite population simulations presented in Section IV. As already mentioned, these equations are derived as the infinite population limit of the imitation dynamics described in Section II (see Fontanari_2024b ). Since $x+y+z=1$ and the expected payoffs of the different strategies depend only on $x$ and $z$, henceforth we will eliminate $y$ in favor of these variables. The resulting equations become
$\frac{dx}{dt}$ $=$ $\frac{x}{\Delta f_{\max}}\left[(1-x)(\pi^{C}-\pi^{D})-z(\pi^{P}-\pi^{D})\right]$ (16)
$\frac{dz}{dt}$ $=$ $\frac{z}{\Delta f_{\max}}\left[(1-z)(\pi^{P}-\pi^{D})-x(\pi^{C}-\pi^{D})\right]$ (17)
where we have omitted the dependence of the expected payoffs of the strategies on $x$ and $z$ for simplicity. Since only payoff differences appear in these equations, we can drop the common term $rc(x+z)$ in Eqs. (6), (7) and (8), which greatly facilitates our analysis of the equilibrium solutions. On the one hand, as $z$ increases from 00 to $1$, $\tilde{\pi}^{C}(z)=\pi^{C}(x,z)-rc(x+z)$ monotonically decreases from $-c$ to $-c-\alpha\beta$. Similarly, $\tilde{\pi}^{D}(z)=\pi^{D}(x,z)-rc(x+z)$ monotonically decreases from 00 to $-\beta$. On the other hand, $\tilde{\pi}^{P}(z)=\pi^{P}(x,z)-rc(x+z)$ monotonically increases from $-c-\gamma$ to $-c-\gamma\zeta$, as $z$ increases from 00 to $1$. Note that the introduction of these shifted payoffs makes it clear that the equilibrium strategy frequencies do not depend on the amplification factor $r$. In addition, by properly rescaling $t$ we can set $c=1$ in the numerical analysis without loss of generality.
The fixed-point solutions of the replicator equations (16) and (17) are obtained by setting $dx/dt=dz/dt=0$. In the following we briefly present these solutions together with the results of a standard local stability analysis Britton_2003 ; Murray_2007 .
III.1 All-cooperators solution
@@ III.2 All-defectors solution
which are always negative and so a population of defectors cannot be invaded, i.e., the all-defectors solution is an evolutionarily stable equilibrium Maynard_1982 . However, the population mean payoff $\bar{\pi}=0$ is the worst possible payoff for a population in equilibrium, which reveals the origin of the public goods dilemma: although the highest individual payoff is obtained by a defector in a group of cooperators ($rc$), if every player decides to defect, the population as a whole ends up in the worst possible situation.
III.3 All-punishers solution
@@ III.4 No-cooperators solution
Figure 1: Unstable fixed point $z^{*}$ corresponding to the no-cooperators solution as a function of the maximum per capita contribution to the sanctioning institution $\gamma$. Left panel: $N=50$ and (from bottom to top) $\zeta=0.1,0.2,0.4$ and $0.8$. Right panel: $\zeta=0.2$ and (from top to bottom at $\gamma=2$) $N=50,100,200$ and $400$. The thick curves are the results for finite $N$, while the thin lines are the results for $N\to\infty$, which exist only for $z^{*}\geq\zeta$. The other parameters are $\beta=3$ and $c=1$.
@@
provided that $\gamma\geq\beta-c$, which ensures that $z^{*}\geq\zeta$ and so that the offenders are punished. For $\gamma<\beta-c$ there is no root for $g(z)$, which means that the size of the domain of attraction of the all-punishers solution vanishes. As expected, Fig. 1 shows the strong finite $N$ effects near the discontinuity predicted by the analytical solution in the limit $N\to\infty$ as well as for $\gamma<\beta-c$.
III.5 No-defectors solution
@@
provided that $\gamma\geq\beta\alpha$ which ensures that $z^{*}\geq\zeta$. Of course, the results for both the finite and infinite play group sizes depend on the product $\alpha\beta$ and not on the individual values of the parameters $\alpha$ and $\beta$. In addition, the parameter $c$ plays no role in the determination of $z^{*}$ for the no-defectors solution. Since the dependence of $z^{*}$ on $\gamma$ is very similar to that shown in Fig. 1, here we look at how the finite play group solutions approach the discontinuous result predicted by the analytical solution in the limit $N\to\infty$. The left panel of Fig.  2 shows the unstable fixed point $z^{*}$ for fixed $\zeta=0.2$ and $\alpha\beta=4$, and a variety of large play group sizes $N$. The right panel shows that the deviation from the $N\to\infty$ prediction vanishes like $1/\sqrt{N}$ as $N$ increases. This is expected since for large $N$ we have used the approximation
$\sum_{K=0}^{K_{m}-1}\binom{N}{K}z^{K}(1-z)^{N-K}$$\approx$ $\frac{1}{2}\mbox{erf}\left[\frac{(\zeta-z)\sqrt{N}}{\sqrt{2z(1-z)}}\right]+$ (26)
$\frac{1}{2}\mbox{erf}\left[\frac{z\sqrt{N}}{\sqrt{2z(1-z)}}\right]$
to derive Eqs. (24) and (25). We note that finding how the deviations scale with $N$ by ‘collapsing’ the curves for different values of $N$ into a single curve is a technique widely used in finite-size scaling analysis to calculate the critical exponents of phase transitions Binder_1985 ; Privman_1990 ; Campos_1999 .
Figure 2: The no-defectors unstable fixed point $z^{*}$ as a function of the maximum per capita contribution to the sanctioning institution $\gamma$ for $\zeta=0.2$, $\alpha\beta=4$ and (from top to bottom at $\gamma=3$) $N=3200,6400,12800$ and $25600$ (left panel). The black thin line is the result for $N\to\infty$. The right panel shows the same curves with the y-axis rescaled to $(z^{*}-\zeta)N^{1/2}$.
Figure 3: Values of $\alpha$ and $\gamma$ (left panel) for which the coexistence fixed point $z^{*}$ (right panel) exists for $\zeta=0.2$, and (from top to bottom at $\alpha=0.6$) $N=50,200,1600$ and $3200$. The other parameters are $\beta=3$ and $c=1$. The vertical dashed lines indicate the threshold $\alpha=1-c/\beta=2/3$ beyond which there is no coexistence solution. The black thin curves indicate the results for $N\to\infty$.
III.6 Coexistence solution
@@
The derivation of the results for $N\to\infty$ is a bit more complicated than in the previous cases. The reason is that as $N$ increases, $z^{*}$ approaches $\zeta$ so that $(\zeta-z^{*})N^{1/2}=\eta$ is finite and nonzero. Using the approximation (26) we can write the equations $g(z^{*})=h(z^{*})=0$ as
@@
For fixed $\beta$, the largest value of $\alpha$ for which Eq. (27) admits a solution is $\alpha=1-c/\beta$, corresponding to $\eta\to-\infty$, and in this case we have $\gamma=\beta-c$. The left panel of Fig. 3 shows the very good agreement between Eq. ( 28) and the numerical results for large $N$. Of course, whenever Eq. (27) admits a solution we have $z^{*}=\zeta$, which is indeed the large $N$ limit of the numerical results, as shown in the right panel of Fig. 3.
This completes the analysis of the fixed points of the replicator equations (16) and (17). Note that there is no fixed point corresponding to the no-punishers scenario (i.e., $z=0$ and $y=1-x$ with $x\in(0,1)$) since $\tilde{\pi}^{D}(0)>\tilde{\pi}^{C}(0)$ for all parameter settings.
III.7 Dynamics
The all-defectors fixed point ($y=1$) is stable for all values of the model parameters, while the all-punishers fixed point ($z=1$) is stable if the conditions $\gamma\zeta<\alpha\beta$ and $\gamma\zeta<\beta-c$ are satisfied. Except for the coexistence and for the no-cooperators fixed points, which are unstable (see figures below), all other fixed points and especially the all-cooperators fixed point ($x=1$) are saddle points. Thus, the useful information we can obtain from the study of the replicator dynamics, i.e., the numerical solution of Eqs. (16) and (17), is the domains of attraction of the stable fixed points, which are better visualized by ternary or simplex plots. In these plots, the vertices of the triangle denote homogeneous populations of punishers (P), cooperators (C), and defectors (D).
@@ III.7 Dynamics
Figure 4: Replicator dynamics for (left panel) $\alpha=0.549$ and $\gamma=1.499$ and (right panel) $\alpha=0.2$ and $\gamma=1.499$. The other parameters are $N=50$, $K_{m}=10$, $\beta=3$ and $c=1$. The filled circles indicate the fixed points. The blue orbits converge to the all-punishers fixed point, the red orbits converge to the all-defectors fixed point, and the green orbit converges to the all-cooperators fixed point. The magenta horizontal line in the left panel is a line of (unstable) coexistence fixed points and is the separatrix delimiting the domains of attraction of P and D. The magenta curve in the right panel is the separatrix over which the dynamics goes to the no-defectors fixed point. The black dashed line is $z=\zeta=0.2$.
@@
Figure 5: Replicator dynamics for (left panel) $\alpha=0.05$ and $\gamma=1.5$ and (right panel) $\alpha=0.9$ and $\gamma=12.5$. The other parameters are $N=50$, $K_{m}=10$, $\beta=3$ and $c=1$. The filled circles indicate the fixed points. The blue orbit converges to the all-punishers fixed point, the red orbits converge to the all-defectors fixed point, and the green orbits converge to the all-cooperators fixed point. The black dashed line is $z=\zeta=0.2$.
@@
Figure 6: Separatrices delimiting the domains of attraction of P and D for $\gamma=1.5$ and (left panel from top to bottom) $\alpha=0.11,012,0.15,0.2,0.3,1$ and $\beta=3$, and (right panel from top to bottom) $\beta=1.6,2,3,4,5,10$ and $\alpha=0.2$. The other parameters are $N=50$, $K_{m}=10$, and $c=1$. The filled circles indicate the fixed points. The black dashed line is $z=\zeta=0.2$. The flow over a separatrix is always towards the no-defectors fixed point.
@@
IV Finite population simulations
As already pointed out, a major advance in the theoretical study of public goods games is the realization that the replicator equation formalism originally introduced to study biological evolution in continuous time Hofbauer_1998 also describes the social dynamics scenario where individuals are more likely to imitate the behavior of their more successful peers Fontanari_2024b . In this section, we offer a comparison between the predictions of the deterministic replicator equations (16) and (17) and the finite population simulations using the stochastic imitation dynamics presented in Section II.

### v3

@@ 3 The replicator equation formulation
$\frac{dx}{dt}$ $=$ $x\left[\pi^{C}(x,z)-\bar{\pi}(x,y,z)\right]\frac{1}{\Delta f_{\max}}$ (13)
$\frac{dy}{dt}$ $=$ $y\left[\pi^{D}(x,z)-\bar{\pi}(x,y,z)\right]\frac{1}{\Delta f_{\max}}$ (14)
$\frac{dz}{dt}$ $=$ $z\left[\pi^{P}(x,z)-\bar{\pi}(x,y,z)\right]\frac{1}{\Delta f_{\max}},$ (15)
where $\bar{\pi}=x\pi^{C}+y\pi^{D}+z\pi^{P}$ is the average payoff of the population and $\Delta f_{\max}$ is given by Eq. (5). Although we could eliminate the factor $1/\Delta f_{\max}$ by a proper rescaling of $t$ and thus obtain the standard form of the replicator equation, keeping it facilitates comparison with the finite population simulations presented in Section 4. As already mentioned, these equations are derived as the infinite population limit of the imitation dynamics described in Section 2 (see [30]). Since $x+y+z=1$ and the expected payoffs of the different strategies depend only on $x$ and $z$, henceforth we will eliminate $y$ in favor of these variables. The resulting equations become
$\frac{dx}{dt}$ $=$ $x\left[(1-x)(\pi^{C}-\pi^{D})-z(\pi^{P}-\pi^{D})\right]\frac{1}{\Delta f_{\max}}$ (16)
$\frac{dz}{dt}$ $=$ $z\left[(1-z)(\pi^{P}-\pi^{D})-x(\pi^{C}-\pi^{D})\right]\frac{1}{\Delta f_{\max}},$ (17)
where we have omitted the dependence of the expected payoffs of the strategies on $x$ and $z$ for simplicity. Since only payoff differences appear in these equations, we can drop the common term $rc(x+z)$ in Eqs. (6), (7) and (8), which greatly facilitates our analysis of the equilibrium solutions. On the one hand, as $z$ increases from 00 to $1$, $\tilde{\pi}^{C}(z)=\pi^{C}(x,z)-rc(x+z)$ monotonically decreases from $-c$ to $-c-\alpha\beta$. Similarly, $\tilde{\pi}^{D}(z)=\pi^{D}(x,z)-rc(x+z)$ monotonically decreases from 00 to $-\beta$. On the other hand, $\tilde{\pi}^{P}(z)=\pi^{P}(x,z)-rc(x+z)$ monotonically increases from $-c-\gamma$ to $-c-\gamma\zeta$, as $z$ increases from 00 to $1$. The introduction of these shifted payoffs makes it clear that the equilibrium strategy frequencies do not depend on the amplification factor $r$.
Since the expected payoffs $\pi^{C}$, $\pi^{D}$, $\pi^{P}$, and the maximum instantaneous payoff difference $\Delta f_{\max}$ are measured in the same payoff units, the time $t$ in the replicator equations (16) and (17) is dimensionless. This is expected because time in the imitation stochastic dynamics is measured in Monte Carlo steps, where a Monte Carlo step corresponds to the choice of $M$ random focal individuals for the update. Finally, we note that by measuring all parameters in units of the contribution $c$ to the public goods, we can set $c=1$ in the numerical analysis without loss of generality.
The fixed-point or equilibrium solutions of the replicator equations (16) and (17), denoted by $x^{*}$ and $z^{*}$, are obtained by setting $dx/dt=dz/dt=0$. In the following we briefly present these solutions together with the results of a standard local stability analysis. This analysis is done by linearizing the replicator equations at the equilibrium solutions, resulting in the linear system
$\begin{pmatrix}\frac{du}{dt}\\ \frac{dv}{dt}\end{pmatrix}=\mathbf{A}\begin{pmatrix}u\\ v\end{pmatrix},$ (18)
where $u=x-x^{*}$ and $v=z-z^{*}$. The $2\times 2$ Jacobian matrix $\mathbf{A}$ evaluated at the equilibrium solutions is called the community matrix and the signs of its eigenvalues determine the local stability of these solutions [47, 48].
3.1 All-cooperators solution
@@ 3.2 All-defectors solution
which are always negative and so a population of defectors cannot be invaded, i.e., the all-defectors solution is an evolutionarily stable equilibrium [5]. However, the population mean payoff $\bar{\pi}=0$ is the worst possible payoff for a population in equilibrium, which reveals the origin of the public goods dilemma: although the highest individual payoff is obtained by a defector in a group of cooperators ($rc$), if every player decides to defect, the population as a whole ends up in the worst possible situation.
3.3 All-punishers solution
@@ 3.4 No-cooperators solution
Figure 1: Unstable fixed point $z^{*}$ corresponding to the no-cooperators solution as a function of the maximum per capita contribution to the sanctioning institution $\gamma$. Left panel: $N=50$ and $\zeta=0.1,0.2,0.4,0.8$, as indicated. Right panel: $\zeta=0.2$ and $N=50,100,200,400$, as indicated. The thick curves are the results for finite $N$, while the thin lines are the results for $N\to\infty$, which exist only for $z^{*}\geq\zeta$. The other parameters are $\beta=3$ and $c=1$.
@@
provided that $\gamma>\beta-c$, which ensures that $z^{*}>\zeta$ and so that the offenders are punished. For $\gamma<\beta-c$ there is no root for $g(z)$, which means that the size of the domain of attraction of the all-punishers solution vanishes. For $z^{*}=\zeta$, equating Eqs. (11) and (12) yields $\gamma=\beta/2-c$. Interestingly, the curves for different $N$ in the right panel of Fig. 1 intersect at the coordinates $\gamma=\beta/2-c=0.5$ and $z^{*}=\zeta=0.2$, signaling a threshold (i.e., a discontinuity) at $z^{*}=\zeta$ when $\gamma$ is considered as a function of $z^{*}$, rather than the reverse. The results show, as expected, the strong effect of demographic noise near the threshold. It is easy to understand the dependence on $N$ shown in the right panel of Fig. 1. On the one hand, for $z^{*}<\zeta$ the fluctuations in the composition of small groups allow a fair probability that $K\geq K_{m}$ in some groups (and so the defectors are punished), guaranteeing the existence of the no-cooperators solution in the region $z^{*}=\zeta$ for finite $N$. For a fixed $\gamma$, the smaller the group size, the lower the value of $z^{*}$ needed to guarantee the existence of this solution, as shown in the figure. On the other hand, if $z^{*}>\zeta$, the same fluctuations allow defectors to go unpunished in some small groups, so for a fixed $\gamma$, the smaller the group size, the higher the value of $z^{*}$ needed to guarantee the existence of the no-cooperators solution. This is the reason why the effect of increasing $N$ is reversed depending on whether $z^{*}$ is greater or less than $\zeta$. In addition, variation of $N$ has no effect at all if $z^{*}=\zeta$.
3.5 No-defectors solution
@@
provided that $\gamma>\beta\alpha$ which ensures that $z^{*}>\zeta$. For $z^{*}=\zeta$ we have $\gamma=\alpha\beta/2$. Of course, the results for both the finite and infinite play group sizes depend on the product $\alpha\beta$ and not on the individual values of the parameters $\alpha$ and $\beta$. In addition, the parameter $c$ plays no role in the determination of $z^{*}$ for the no-defectors solution. Since the dependence of $z^{*}$ on $\gamma$ is very similar to that shown in Fig. 1, here we look at how the finite play group solutions approach the discontinuous result predicted by the analytical solution in the limit $N\to\infty$. The left panel of Fig.  2 shows the unstable fixed point $z^{*}$ for fixed $\zeta=0.2$ and $\alpha\beta=4$, and a variety of large play group sizes $N$. The right panel shows that the deviation from the $N\to\infty$ prediction vanishes like $1/\sqrt{N}$ as $N$ increases. This is expected since for large $N$ we have used the approximation (see A)
$\sum_{K=0}^{K_{m}-1}\binom{N}{K}z^{K}(1-z)^{N-K}\approx\frac{1}{2}\mbox{erf}\left[\frac{(\zeta-z)\sqrt{N}}{\sqrt{2z(1-z)}}\right]+\frac{1}{2}\mbox{erf}\left[\frac{z\sqrt{N}}{\sqrt{2z(1-z)}}\right]$ (27)
to derive Eqs. (25) and (26). We note that finding how the deviations scale with $N$ by ‘collapsing’ the curves for different values of $N$ into a single curve is a technique widely used in finite-size scaling analysis to calculate the critical exponents of phase transitions (threshold phenomena) [33, 34, 35]. In fact, there is a threshold at $z^{*}=\zeta$, indicated by the crossing of the curves for different $N$ at $\gamma=\alpha\beta/2$ and $z^{*}=\zeta$. As before, the effect of increasing $N$ is reversed depending on whether $z^{*}$ is greater or less than $\zeta$. For $N\to\infty$, the punishers can be maintained in the population only if $z^{*}>\zeta$, but for finite $N$ there is a fair probability that $K\geq K_{m}$ in some groups even for $z^{*}<\zeta$, which is enough to maintain the punishers in the population.
Figure 2: The no-defectors unstable fixed point $z^{*}$ as a function of the maximum per capita contribution to the sanctioning institution $\gamma$ for $\zeta=0.2$, $\alpha\beta=4$, and $N=3200,6400,12800,25600$, as indicated (left panel). The black thin line is the result for $N\to\infty$. The right panel shows the same curves with the y-axis rescaled to $(z^{*}-\zeta)N^{1/2}$. The color convention is the same for both panels.
Figure 3: Values of $\alpha$ and $\gamma$ (left panel) for which the coexistence fixed point $z^{*}$ (right panel) exists for $\zeta=0.2$, and $N=50,200,1600,3200$, as indicated. The color convention is the same for both panels. The other parameters are $\beta=3$ and $c=1$. The vertical dashed lines indicate the threshold $\alpha=1-c/\beta=2/3$ beyond which there is no coexistence solution. The black thin curves indicate the results for $N\to\infty$.
3.6 Coexistence solution
@@
The derivation of the results for $N\to\infty$ is a bit more complicated than in the previous cases. The reason is that as $N$ increases, $z^{*}$ approaches $\zeta$ so that $(\zeta-z^{*})N^{1/2}=\eta$ is finite and nonzero (see A). Using the approximation (27) we can write the equations $g(z^{*})=h(z^{*})=0$ as
@@
For fixed $\beta$, the largest value of $\alpha$ for which Eq. (28) admits a solution is $\alpha=1-c/\beta$, corresponding to $\eta\to-\infty$, and in this case we have $\gamma=\beta-c$. The left panel of Fig. 3 shows the very good agreement between Eq. ( 29) and the numerical results for large $N$. Of course, whenever Eq. (28) admits a solution we have $z^{*}=\zeta$, which is indeed the large $N$ limit of the numerical results, as shown in the right panel of Fig. 3. This panel shows that for finite $N$, somewhat counterintuitively, increasing $\alpha$ does not favor the punishers, since $z^{*}$ increases with $\alpha$, resulting in a reduction of the basin of attraction of the all-punishers solution. The reason is that for the coexistence solution to exist, $\gamma$ must increase with $\alpha$ (see the left panel of Fig. 3), which is detrimental to the punishers.
This completes the analysis of the fixed points of the replicator equations (16) and (17). There is no fixed point corresponding to the no-punishers scenario (i.e., $z=0$ and $y=1-x$ with $x\in(0,1)$) since $\tilde{\pi}^{D}(0)>\tilde{\pi}^{C}(0)$ for all parameter settings.We note that since the eigenvalues of the community matrices for the stable (i.e., all-punishers and all-defectors) fixed points do not depend on $N$, the group size only affects the sizes of the domains of attraction of these fixed points.
3.7 Dynamics
The all-defectors fixed point ($y=1$) is stable for all values of the model parameters, while the all-punishers fixed point ($z=1$) is stable if the conditions $\gamma\zeta<\alpha\beta$ and $\gamma\zeta<\beta-c$ are satisfied. Except for the coexistence and for the no-cooperators fixed points, which are unstable (see figures below), all other fixed points and especially the all-cooperators fixed point ($x=1$) are saddle points. Thus, the useful information we can obtain from the study of the replicator dynamics, i.e., the numerical solution of Eqs. (16) and (17), is the domains of attraction of the stable fixed points, which are better visualized by ternary or simplex plots. In these plots, the vertices of the triangle denote homogeneous populations of punishers (P), cooperators (C), and defectors (D). As usual, we represent the stable points by a filled (black) circle and an unstable or a saddle point by an open (white) circle.
@@ 3.7 Dynamics
Figure 4: Replicator dynamics for (left panel) $\alpha=0.549$ and $\gamma=1.499$ and (right panel) $\alpha=0.2$ and $\gamma=1.499$. The other parameters are $N=50$, $K_{m}=10$, $\beta=3$ and $c=1$. The filled circles indicate the stable fixed points and the open circles the saddle and unstable fixed points. The blue orbits converge to the all-punishers fixed point, the red orbits converge to the all-defectors fixed point, and the green orbits converge to the all-cooperators fixed point as indicated by the arrows in representative orbits. The magenta horizontal line in the left panel is a line of (unstable) coexistence fixed points and is the separatrix delimiting the domains of attraction of P and D. The magenta curve in the right panel is the separatrix over which the dynamics goes to the no-defectors fixed point, as indicated by the arrow. The black dashed line is $z=\zeta=0.2$.
@@
Figure 5: Replicator dynamics for (left panel) $\alpha=0.05$ and $\gamma=1.5$ and (right panel) $\alpha=0.9$ and $\gamma=12.5$. The other parameters are $N=50$, $K_{m}=10$, $\beta=3$ and $c=1$. The filled circles indicate the stable fixed points and the open circles the saddle and unstable fixed points. The blue orbits converge to the all-punishers fixed point, the red orbits converge to the all-defectors fixed point, and the green orbits converge to the all-cooperators fixed point as indicated by the arrows in representative orbits. The black dashed line is $z=\zeta=0.2$.
@@
Figure 6: Separatrices delimiting the domains of attraction of P and D for $\gamma=1.5$ and (left panel from top to bottom) $\alpha=0.11,012,0.15,0.2,0.3,1$ and $\beta=3$, and (right panel from top to bottom) $\beta=1.6,2,3,4,5,10$ and $\alpha=0.2$. The other parameters are $N=50$, $K_{m}=10$, and $c=1$. The filled circles indicate the stable fixed points and the open circles the saddle and unstable fixed points. The flow over a separatrix is always towards the no-defectors fixed point, as indicated by the arrow in representative orbits. The black dashed line is $z=\zeta=0.2$.
@@
A word is in order about the effect of changing the group size $N$ on the orbits shown in the above simplexes. As noted earlier, changing $N$ only affects the sizes of the domains of attraction of the stable fixed points, so its effect is negligible in the scenarios shown in Fig. 5, where the only stable fixed point is the all-defectors fixed point. In the more interesting scenario of bistability between the all-defectors and all-punishers fixed points shown in Fig. 4, increasing $N$ simply shifts the separatrices slightly. For example, considering the parameter setting used in the right panel of Fig. 4, for $N=10$ and $K_{m}=2$ we find that the equilibrium frequency of punishers for the no-defectors fixed point is $z_{nd}^{*}=0.506$, while for the no-cooperators fixed point it is $z_{nc}^{*}=0.242$. We recall that these fixed points determine the end points of the separatrix. For $N\to\infty$ and $\zeta=0.2$, we find $z_{nd}^{*}=0.5$ and $z_{nc}^{*}=0.2$, so the separatrices for $N=10$ and $N\to\infty$ differ very little.
4 Finite population simulations
As already pointed out, a major advance in the theoretical study of public goods games is the realization that the replicator equation formalism originally introduced to study biological evolution in continuous time [29] also describes the social dynamics scenario where individuals are more likely to imitate the behavior of their more successful peers [30]. In this section, we offer a comparison between the predictions of the deterministic replicator equations (16) and (17) and the finite population simulations using the stochastic imitation dynamics presented in Section 2.